# Should we pre-train a decoder in contrastive learning for dense prediction tasks?

### v1

@@ Abstract
Contrastive learning in self-supervised settings primarily focuses on pre-training encoders, while decoders are typically introduced and trained separately for downstream dense prediction tasks. This conventional approach, however, overlooks the potential benefits of jointly pre-training both the encoder and decoder. In this paper, we propose DeCon: a framework-agnostic adaptation to convert an encoder-only self-supervised learning (SSL) contrastive approach to an efficient encoder-decoder framework that can be pre-trained in a contrastive manner. We first update the existing architecture to accommodate a decoder and its respective contrastive loss. We then introduce a weighted encoder-decoder contrastive loss with non-competing objectives that facilitates the joint encoder-decoder architecture pre-training. We adapt two established contrastive SSL frameworks tailored for dense prediction tasks, achieve new state-of-the-art results in COCO object detection and instance segmentation, and match state-of-the-art performance on Pascal VOC semantic segmentation. We show that our approach allows for pre-training a decoder and enhances the representation power of the encoder and its performance in dense prediction tasks. This benefit holds across heterogeneous decoder architectures between pre-training and fine-tuning and persists in out-of-domain, limited-data scenarios.
@@ 1 Introduction
The increasing demand for efficient Deep Learning (DL) methods arises from their ability to address complex tasks. While traditional approaches like supervised learning can yield powerful, high-performing models, they depend heavily on large volumes of high-quality labeled data. However, obtaining such annotations is often tedious, expensive, and sometimes impractical, even when abundant data is available. To overcome this bottleneck, one can benefit from a pre-trained model. Pre-trained models are usually models trained from scratch on a large scale dataset and allow for better performance on downstream tasks with smaller datasets. Typically, an encoder is pre-trained to capture a representation of the input data, which can then be fine-tuned for downstream tasks such as classification or segmentation. During the downstream fine-tuning phase, randomly initialized layers are typically added on top of the encoder, and the complete architecture is retrained using a supervised learning approach on part of or on a full annotated dataset. ImageNet [15] is a large and widely utilized dataset with annotated classes. However, models pre-trained on ImageNet classification are not the best fit for every downstream tasks, especially when the downstream objective is not classification. Since creating large-scale annotated dataset for every tasks presents numerous challenges, Self-Supervised Learning (SSL) has emerged as a promising alternative, enabling models to learn from the vast amount of unlabeled data and perform effectively on various downstream tasks. These pre-trained models can reach downstream performance close to a fully supervised approach using much smaller to no labeled datasets. Notable approaches such as SimCLR [7], VicReg [2], and MoCo [22] showed promising results by leveraging contrastive learning for downstream classification. However, most SSL frameworks primarily target classification downstream tasks. Some efforts such as PixCon [37] or DenseCL [45], focus on dense prediction tasks like segmentation or object detection. In all of these efforts, authors only train the encoders of the DL architectures and focus on incorporating a local loss with existing classification-targeted SSL approaches to adapt with dense tasks.
@@
In this work, we propose a Decoder-aware contrastive learning (DeCon) approach, a framework agnostic encoder-decoder adaptation to pre-train an encoder jointly with a decoder, improving its representation power and efficiently preparing it to be combined with a decoder in downstream tasks. Our contributions are as follows.
@@
  • We propose DeCon-SL, a single-level joint encoder-decoder adaptation of contrastive dense SSL frameworks. We show that jointly pre-training the encoder with a decoder enhances the representation power of the encoder and improves downstream dense prediction performance, even without transferring the pre-trained decoder.
@@
  • We improve on this adaptation by introducing DeCon-ML, which extends DeCon-SL with a multi-level decoder loss function and an encoder level-wise channel dropout to promote a comprehensive feature usage of the encoder. It achieves state-of-the-art results on COCO object detection and instance segmentation and matches top performance on Pascal VOC semantic segmentation.
@@
  • We demonstrate that DeCon improves the encoder’s downstream performance compared to encoder-only frameworks, even when pre-training and fine-tuning decoders differ.
@@
  • We show that DeCon’s efficiency persists in out-of-domain downstream tasks and provides even greater benefits in scenarios with limited annotated fine-tuning data.
@@ 2 Related works
Self-Supervised pre-training: In SSL, deep learning models are trained on pretext tasks where the data itself provides supervision, which allow models to learn meaningful representations without any manual annotations. Early SSL works focused on designing these tasks, such as predicting missing parts of an image (image inpainting) [38], context prediction [16], solving jigsaw puzzles [34], colorizing grayscale images [52], and predicting rotations [20]. These methods aimed to create surrogate objectives that encourage models to develop feature representations, which could later be transferred to downstream tasks. However, these methods often produced task-specific representations with limited generalizability and restricted scalability.
@@
Recent methods prioritize generalizability and move away from hand-crafted pretext tasks. These approaches can be broadly categorized into generative and contrastive methods. Generative approaches in SSL aim to learn meaningful representations by modeling the underlying data distribution. This is typically achieved through reconstruction-based methods, where the model predicts corrupted parts of the input [23, 1, 17, 51], or adversarial methods, where a generative model learns to synthesize realistic samples while an encoder captures high-level features [48, 26]. Contrastive methods, which are the focus of this research, learn data representations by maximizing the similarity between features extracted from two augmented views of the same image. To avoid model collapse—a phenomenon where the model maps all inputs to the same representation—negative pairs [22, 8], stop-gradient mechanisms [9], variance regularization [2, 3], momentum contrast [22] or other architectural and optimization techniques are employed [5].
@@
Dense pre-training: Even though most of the aforementioned frameworks can be used for downstream segmentation or detection, these models are not designed specifically for pixel-level, i.e., dense downstream tasks. Consequently, developing SSL approaches that generalize well to dense downstream tasks, such as the segmentation task, is an active area of research [46, 45, 44, 3, 35]. The main idea behind these approaches is to shift from a global image similarity to a local similarity. Local similarity can be enforced based on the location of pixels in an image [50, 44, 35], local pixel features [45], or both [29, 3, 37]. Local similarity can also be enforced in a region-based manner [19, 24, 49, 46]. These methods outperform those based solely on image-level global similarity in downstream dense tasks.
@@
Encoder-Decoder pre-training: Encoder-decoder architectures with skip connections between the encoder and decoder layers (e.g., U-Net-like models) have achieved state-of-the-art results in many segmentation tasks [40, 27, 53]. A common approach for pre-training such models in a self-supervised manner is image reconstruction [30, 4, 6]. However, the presence of skip connections allows a substantial amount of information to bypass the encoder’s final layer—often referred to as the bottleneck layer–resulting in lower-quality embeddings generated by the encoder. Our proposed approach preserves skip connections while addressing the issue of inferior representations generated by the decoder. Encoder-only contrastive learning frameworks are built solely around an encoder, incorporating components such as predictors and projectors. However, they do not include a decoder, even though a randomly initialized decoder is later attached to the pre-trained encoder in downstream dense tasks to form a complete architecture. While some generative SSL frameworks use a decoder for pre-training [14, 38, 23], to the best of our knowledge, no work has explored pre-training vision models with both encoder and decoder loss in a contrastive learning framework to learn a data representation.
@@
Figure 1: A: DeCon-SL. Instead of a classical encoder-only pre-training, a decoder is pre-trained alongside the encoder. We mirror the encoder loss at the decoder level and optimize the architecture using a weighted sum of the encoder and decoder losses. B: DeCon-ML. Instead of computing the decoder loss at a single level, it is calculated across multiple levels (four in this figure). Additionally, a channel-wise dropout is applied at the output of each encoder level before it is passed through the skip connection to the decoder.
@@ 3.1 DeCon-SL
The DeCon-SL architecture is illustrated in Fig. 1 A, where $f_{\theta}$ and $f_{\epsilon}$ represent teacher and student encoders, respectively. $aux^{enc}_{\theta}$ and $aux^{enc}_{\epsilon}$ are the SSL framework’s auxiliary layers of the teacher and student encoders, respectively. Auxiliary layers can include any projector, predictor heads, or any layers used in the loss definition of the original SSL framework. $g_{\theta}$ and $g_{\epsilon}$ are the decoders of the teacher and student networks, and $aux^{dec}_{\theta}$ and $aux^{dec}_{\epsilon}$ are the auxiliary layers feeding from their respective decoder. $g_{\theta}$ and $g_{\epsilon}$ take the output of encoder $f_{\theta}$ and $f_{\epsilon}$ as input, respectively. Auxiliary layers $aux^{enc}_{\theta}$ and $aux^{enc}_{\epsilon}$ are responsible for generating a representation from the encoder output, that is used to compute a contrastive loss for the encoders $L^{encoder}$. The contrastive loss, $L^{decoder}$ is calculated from the representation obtained from the auxiliary layers of the decoders, i.e., $aux^{dec}_{\theta}$ and $aux^{dec}_{\epsilon}$. $f_{\epsilon}$ and $g_{\epsilon}$ are updated using the exponential moving average (EMA) of $f_{\theta}$ and $g_{\theta}$, respectively.
@@ 3.2 DeCon-ML
To address the challenge posed by skip connections between the encoder and decoder layers (as mentioned in the Related works section), we introduce a multi-layer contrastive decoder loss, consisting of a single loss for the encoder and multiple loss components for the decoder along with channel dropout and deep supervision, which we describe in detail below. Figure 1 B illustrates the DeCon-ML adaptation of a given SSL framework.
Channel dropout: The dropout layer is typically used to prevent overfitting, improve generalization, and act as an implicit ensemble [41, 43]. However, we extend its use by applying dropout to channels transferred through skip connections between the encoder and decoder. This prevents the model from relying heavily on specific features shared through skip connections and enforces full encoder training, leading to richer data representations at the bottleneck layer of the encoder. In DeCon-ML, we apply channel dropout to the output layers of the encoder at the different levels when passing these outputs to the decoder. No channel is zeroed-out when the inputs pass through the encoder. This channel dropout promotes a more comprehensive utilization of the encoder parameters and enhances representation power at various encoder levels while preserving the input information content as it passes through the encoder to its bottleneck.
@@ 3.3 Implementation details
Architecture: Following common practice in contrastive SSL approaches literature, we use a ResNet-50 encoder [21] as a backbone, a Feature Pyramid Network (FPN) [32] for downstream detection and instance segmentation, and a Fully Convolutional Network (FCN) [33] for downstream semantic segmentation [45, 22, 46] tasks. We implement FCN and skip-connected FPN as two decoder architectures in the DeCon framework adaptation. We follow the mmseg [11] FCN implementation used in SlotCon [46] and customize FPN implementation from Detectron2 [47] version 0.6. We use DeCon to adapt the SlotCon [46] and DenseCL [45] frameworks. More implementation details are available in Supplementary Material Section 1.
@@
For SlotCon, the student and teacher encoders, and every auxiliary layers—i.e., the respective projectors and semantic groupings, the slot predictor and the prototypes—are kept unchanged. Two decoders, a student and a teacher decoder (either FCN or FPN), that take the output of their corresponding encoder as input are added to the framework. The auxiliary layers are replicated at the decoder level. In the case of decoder deep supervision with FPN decoders, auxiliary layers are replicated independently at the four decoder levels. No deep supervision is performed in the DeCon adaptation of SlotCon with an FCN decoder. The decoder projectors differ from the encoder projectors only in input channels: 2048 for the encoder and 256 for the FPN and FCN decoders. We follow SlotCon by updating the teacher branch and the encoder and decoder slot predictors with Backpropagation, the student branch with an EMA of the first one, and the prototypes centers with an EMA of the previous batch centers seen during training. Our implementation is based on the official SlotCon repository. Detailed figures of SlotCon DeCon-SL and DeCon-ML adaptations are available in the Supplementary Material Section 1.
For DenseCL, the two encoders, along with their respective auxiliary layers—i.e., the global and dense projection heads—are kept the same as well as the independent dictionaries for global and dense losses. Additionally, we introduce two decoders, which take their respective encoders’ outputs as inputs. We also replicate the encoder’s dense and global projection heads and dictionaries separately for the two decoders. As for SlotCon, the input dimension of the decoder projectors is 256 instead of 2048 for the encoder. The encoders and decoders dictionaries size is set to 16384. We follow DenseCL by updating the first branch of the architecture through Backpropagation, while the second branch is updated using an EMA of the first one. Due to extensive memory requirements of replicating dictionaries at each decoder level, we do not conduct decoder deep supervision or channel dropout (DeCon-ML) experiments with DenseCL. We adapt DenseCL implementation from mmselfsup package [12] version 1.0.0.
@@
Pre-training setup: We pre-train DeCon on COCO 2017 dataset [31]. The augmentations and all the hyperparameters—learning rate, weight decay, schedulers, batch size, input shape, number of epochs, momentum, and so on—are kept the same as used in the original SSL framework: SlotCon or DenseCL. The SlotCon models are pre-trained for 800 epochs using a LARS optimizer, with a total batch size of 512 across all available GPUs. The base learning rate is 1.0 scaled with the batch size and updated with a cosine learning rate decay schedule using weight decay of $10^{-5}$ and a warm-up period of 5 epochs. The pre-trainings using FPN decoders are done using eight NVIDIA H100 GPUs (80GB each). The pre-trainings without decoder and with FCN decoders are conducted on one 80GB NVIDIA A100 GPU. The DenseCL models are pre-trained for 800 epochs using an SGD optimizer with a total batch size of 256. The base learning rate is 0.3 and is updated using a cosine learning rate decay schedule (0.0001 weight decay, 0.9 momentum). The pre-training experiments are conducted with eight NVIDIA A100 GPUs (40GB each).
@@
To assess the DeCon framework on non scene-centric data, we pre-train SlotCon and a DeCon-SL adaptation of SlotCon with an FCN decoder on two small-scale datasets: ISIC 2017 and REFUGE described in Sec. 3.4. For ISIC dataset, we pre-train for 800 epochs with a batch size of 256 and a base learning rate of 1.0 linearly scaled with batch size. For REFUGE, we pre-train for 2400 epochs with a batch size of 192 and the same learning rate as for ISIC. These pre-trainings are performed with one NVIDIA A6000 GPU with 48GB of memory.

### v2

# Beyond the Encoder: Joint Encoder-Decoder Contrastive Pre-Training Improves Dense Prediction

@@ Abstract
Contrastive learning methods in self-supervised settings have primarily focused on pre-training encoders, while decoders are typically introduced and trained separately for downstream dense prediction tasks. However, this conventional approach overlooks the potential benefits of jointly pre-training both encoder and decoder. In this paper, we propose DeCon, an efficient encoder-decoder self-supervised learning (SSL) framework that supports joint contrastive pre-training. We first extend existing SSL architectures to accommodate diverse decoders and their corresponding contrastive losses. Then, we introduce a weighted encoder-decoder contrastive loss with non-competing objectives to enable the joint pre-training of encoder-decoder architectures. By adapting an established contrastive SSL framework for dense prediction tasks, DeCon achieves new state-of-the-art results: on COCO object detection and instance segmentation when pre-trained on COCO dataset; across almost all dense downstream benchmark tasks when pre-trained on COCO+ and ImageNet-1K. Our results demonstrate that joint pre-training enhances the representation power of the encoder and improves performance in dense prediction tasks. This gain persists across heterogeneous decoder architectures, various encoder architectures, and in out-of-domain limited-data scenarios.
@@ 1 Introduction
The growing demand for efficient Deep Learning (DL) methods stems from their ability to address complex tasks. While traditional approaches like supervised learning can yield high-performing models, they depend on large volumes of high-quality annotated data. Obtaining such annotations is often labor-intensive, expensive, and sometimes impractical, even when abundant data is available. A common workaround is to use pre-trained models, typically encoders trained from scratch on large-scale datasets, to boost performance on downstream tasks with limited labels. Typically, an encoder is pre-trained to capture a representation of the input data, which can then be fine-tuned for downstream tasks such as classification or segmentation. During the downstream fine-tuning phase, randomly initialized layers are typically added on top of the encoder, and the complete architecture is retrained using a supervised learning approach on an annotated dataset. ImageNet [15] is a widely utilized labeled dataset for such pre-training. However, models pre-trained on ImageNet classification task often transfer poorly to downstream dense prediction tasks, such as detection and segmentation [56, 47, 41].
Creating large-scale annotated datasets for every task is impractical. Self-Supervised Learning (SSL) is a promising alternative, enabling models to learn from the large-scale unannotated data to enhance performance on various downstream tasks. These pre-trained models can reach downstream performance close to a fully supervised approach while using significantly fewer labeled examples. Notable approaches such as SimCLR [7], VicReg [2], and MoCo [22] showed promising results by leveraging contrastive learning for downstream classification. However, most SSL frameworks primarily target classification-based downstream tasks. Some efforts, such as PixCon [39] and DenseCL [47], focus on dense prediction tasks like segmentation or object detection. However, these approaches only pre-train encoders for downstream use and rely on local losses adapted from classification-oriented SSL methods without jointly training the decoder.
@@
In this work, we propose a Decoder-aware contrastive learning (DeCon) approach, an encoder-decoder framework that jointly pre-trains encoder and decoder in an SSL manner. DeCon  enhances the representation power of the encoder and prepares it more effectively for integration with decoders in downstream dense prediction tasks. The contributions of this paper are as follows:
@@
  • We propose DeCon-SL, a single-level joint encoder-decoder adaptation of contrastive dense SSL frameworks. We show that jointly pre-training the encoder and decoder enhances the representation power of the encoder and improves downstream dense prediction performance, even when transferring only the pre-trained decoder.
@@
  • We further extend this adaptation by introducing DeCon-ML, which builds on DeCon-SL with a multi-level decoder loss function and an encoder-level channel dropout to promote a comprehensive feature utilization of the encoder. DeCon-ML achieves new state-of-the-art results in COCO object detection and instance segmentation when pre-trained on COCO dataset; and across almost all dense downstream benchmark tasks when pre-trained on COCO+ and ImageNet-1K.
@@
  • We demonstrate that DeCon improves the encoder performance in downstream tasks compared to encoder-only frameworks, even when the decoders used for pre-training and fine-tuning differ.
  • We show that DeCon can achieve gains without increasing parameter count or GPU requirements relative to the original framework.
@@
  • We demonstrate that DeCon remains effective in out-of-domain downstream tasks and yield even greater benefits in scenarios with limited annotated data for fine-tuning.
@@ 2 Related Works
Self-Supervised pre-training: In early SSL works, deep learning models were trained on pretext tasks where supervision was derived directly from the data itself. This enabled the models to learn meaningful representations without relying on manual annotations. These tasks included predicting missing parts of an image [40], context prediction [16], solving jigsaw puzzles [36], colorizing grayscale images [54], and predicting rotations [20]. These methods aimed to create surrogate objectives that encouraged models to develop feature representations transferable to downstream tasks. However, they often produced task-specific representations with limited generalizability.
@@
Recent SSL methods prioritize generalizability and move away from hand-crafted pretext tasks. These approaches can be broadly categorized into generative and contrastive methods. Generative approaches aim to learn meaningful representations by modeling the underlying data distribution. This is typically achieved through reconstruction-based methods, where the model predicts corrupted parts of the input [23, 1, 17, 53], or through adversarial methods, where a generator model learns to synthesize realistic samples to challenge a discriminator, thereby encouraging the encoder to extract informative semantic features [50, 26]. Contrastive methods, which are the focus of this research, learn data representations by maximizing the similarity between features extracted from two augmented views of the same image. To avoid model collapse—a phenomenon where the model maps all inputs to the same representation—strategies such as negative sampling [22, 8], stop-gradient mechanisms [9], variance regularization [2, 3], momentum encoders [22], and other architectural and optimization techniques have been employed [5].
@@
Dense pre-training: Although most of the aforementioned frameworks can be used for downstream segmentation or detection, they are not designed specifically for pixel-level—i.e., dense prediction downstream tasks such as instance and semantic segmentation. Consequently, developing SSL approaches that generalize well to such downstream tasks remains an active area of research [48, 47, 46, 3, 37]. The main idea behind these approaches is to shift from a global image similarity to a local similarity. They aim to enforce local similarity based on the location of pixels in an image [52, 46, 37], local pixel features [47], or both [29, 3, 39]. Local similarity can also be enforced in a region-based manner [19, 24, 51, 48]. These methods outperform those relying solely on image-level global similarity when evaluated on dense prediction downstream tasks.
@@
Encoder-only versus encoder-decoder pre-training learning: Encoder-decoder architectures with skip connections between the encoder and decoder layers (i.e., the U-Net family of models) have achieved state-of-the-art results in many segmentation tasks [43, 27, 55]. A common approach for pre-training such models in a self-supervised manner is image reconstruction [30, 4, 6]. However, the presence of skip connections allows a substantial amount of information to bypass the encoder’s final layer—often referred to as the bottleneck layer–resulting in lower-quality embeddings generated by the encoder. Our proposed approach preserves skip connections while addressing the issue of inferior representations generated by the decoder.
Encoder-only contrastive learning frameworks are built solely around an encoder, incorporating components such as predictors and projectors. However, they do not include a decoder. After pre-training, a randomly initialized decoder is attached to the pre-trained encoder in dense prediction downstream tasks to form a complete architecture. While some generative SSL frameworks use a decoder for pre-training [14, 40, 23], to the best of our knowledge, no work has explored pre-training vision models with both encoder and decoder loss in a contrastive learning framework to learn informative data representations.
@@
Figure 1: A: DeCon-SL. Instead of a classical encoder-only pre-training, a decoder is pre-trained alongside the encoder. We mirror the encoder loss at the decoder level and optimize the architecture using a weighted sum of encoder and decoder losses. B: DeCon-ML. Instead of computing the decoder loss at a single level, it is calculated across multiple levels (four in this figure). Additionally, a channel-wise dropout is applied at the output of each encoder level before it is passed through the skip connection to the decoder.
@@ 3.1 DeCon-SL
The DeCon-SL architecture is illustrated in Fig. 1 A, where $f_{\theta}$ and $f_{\epsilon}$ represent student and teacher encoders, respectively. $aux^{enc}_{\theta}$ and $aux^{enc}_{\epsilon}$ are the SSL framework’s auxiliary layers of the student and teacher encoders, respectively. Auxiliary layers can include any projector, predictor heads, or any layers used in the loss definition of the original SSL framework. $g_{\theta}$ and $g_{\epsilon}$ are the decoders of the student and teacher networks, and $aux^{dec}_{\theta}$ and $aux^{dec}_{\epsilon}$ their respective auxiliary layers. $g_{\theta}$ and $g_{\epsilon}$ take the output of encoder $f_{\theta}$ and $f_{\epsilon}$ as input, respectively. Auxiliary layers $aux^{enc}_{\theta}$ and $aux^{enc}_{\epsilon}$ are responsible for generating a representation from the encoder output, that is used to compute a contrastive loss for the encoders $L^{encoder}$. The contrastive loss, $L^{decoder}$ is calculated from the representation obtained from the auxiliary layers of the decoders, i.e., $aux^{dec}_{\theta}$ and $aux^{dec}_{\epsilon}$. $f_{\epsilon}$ and $g_{\epsilon}$ are updated using the exponential moving average (EMA) of $f_{\theta}$ and $g_{\theta}$, respectively.
@@ 3.2 DeCon-ML
To promote a comprehensive usage of the encoder’s parameters, we introduce a multi-layer contrastive decoder loss. This updated framework consists of a single loss for the encoder and multiple loss components for the decoder in a deep supervision manner along with channel dropout, which we describe in detail below. Figure 1 B illustrates the DeCon-ML adaptation of a given SSL framework.
Channel dropout: The dropout layer is typically used to prevent overfitting, improve generalization, and act as an implicit ensemble [44, 45]. However, we extend its use by applying dropout to channels transferred through skip connections between the encoder and decoder. In DeCon-ML, we apply channel dropout to the output layers of the different encoder levels by zero-ing out entire channels of the feature maps when passing these outputs to the decoder. No channel is zeroed-out when the inputs pass through the encoder. This channel dropout prevents the model from over-relying on specific features shared through skip connections. It encourages a more comprehensive use of the encoder’s parameters at the different levels, leading to a richer and more powerful data representation learned, while still preserving the input information content as it passes through the encoder to its bottleneck.
@@ 3.3 Implementation details
Architecture: We use a ResNet-50 encoder [21] as a backbone and implement a Fully Convolutional Network (FCN) [34] and a skip-connected Feature Pyramid Network (FPN) [32] as two decoder architectures in the DeCon framework. We follow the mmseg [11] FCN implementation used in SlotCon [48] and customize the FPN implementation from Detectron2 [49] version 0.6. Details on these decoder architectures are provided in Supplementary Material Section 1. We use DeCon to adapt the SlotCon [48] framework. Additional experiments are conducted using DenseCL [47] and PixPro [52] framework or a ConvNeXt-small [33] backbone.
@@
For SlotCon, the student and teacher encoders, and every auxiliary layers—i.e., the respective projectors and semantic groupings, the slot predictor and the prototypes—are kept unchanged. Two decoders, a student and a teacher decoder (either FCN or FPN), that take the output of their corresponding encoder as input are added to the framework. The auxiliary layers are replicated at the decoder level. In the case of decoder deep supervision with FPN decoders (DeCon-ML), auxiliary layers are replicated independently at the four decoder levels. No deep supervision is performed in the DeCon adaptation of SlotCon with an FCN decoder (DeCon-SL). The decoder projectors differ from the encoder projectors only in input channels: 2048 for the encoder and 256 for the FPN and FCN decoders. We follow SlotCon by updating the student branch and the corresponding encoder and decoder slot predictors with Backpropagation, the teacher branch with an EMA of the first one, and the prototypes centers with an EMA of the previous batch centers seen during training. Our implementation is based on the official SlotCon repository.
We adapt in a similar way DenseCL and PixPro into a DeCon-SL framework using an FCN decoder. More details are given in Supplementary Material 2. We adapt DenseCL implementation from mmselfsup package [12] version 1.0.0 and PixPro from the original implementation.
Although the backbone pre-trained remains unchanged, adding multiple decoder levels and associated loss functions increases the overall number of parameters in the DeCon-ML architecture. To enable a fair comparison with the original SlotCon framework using a similar parameter budget, we implemented a reduced version of DeCon-ML that includes only the first two decoder levels originating from the encoder bottleneck. Additionally, the hidden dimension in the decoder projector was reduced from 4096 to 2048. This smaller version, which matches the parameter count of the original SlotCon framework (when $\alpha=0$), is referred to as DeCon-ML-S, while the full model with four decoder levels and unmodified configuration is denoted DeCon-ML-L. Detailed figures of DeCon-SL, DeCon-ML-S and DeCon-ML-L adaptations of SlotCon are available in the Supplementary Material Section 3.
@@
Pre-training setup: We pre-train DeCon on COCO 2017 dataset [31], COCO+ (train2017 + unlabeled2017 datasets) and ImageNet-1K datasets [15]. The augmentations and hyperparameters—learning rate, weight decay, schedulers, input shape, momentum, and so on—are kept the same as used in the original SSL framework. Models are pre-trained on COCO and COCO+ for 800 epochs, and for 200 epochs on ImageNet. The SlotCon ResNet-50 models are pre-trained using a LARS optimizer, with a total batch size of 512 across all available GPUs. The base learning rate is 1.0 scaled with the batch size and updated with a cosine learning rate decay schedule using weight decay of $10^{-5}$ and a warm-up period of 5 epochs. The ConvNeXt-small backbone is pre-trained in SlotCon and DeCon-ML using an AdamW optimizer [35], a base learning rate of 0.002 and weight-decay 0.05. Remaining hyperparameters are kept the same as for the ResNet-50 implementation. The pre-trainings using FPN decoders are done using eight NVIDIA H100 GPUs (80GB each). The pre-trainings without decoder and with FCN decoders are conducted on one 80GB NVIDIA A100 GPU.
For DenseCL, DeCon-SL is pre-trained using an SGD optimizer with a total batch size of 256. The base learning rate is 0.3 and is updated using a cosine learning rate decay schedule (0.0001 weight decay, 0.9 momentum). The pre-training experiments are conducted with eight NVIDIA A100 GPUs (40GB each). For PixPro, DeCon-SL is pre-trained using a LARS optimizer for 800 epochs with a batch size of 512 and a learning rate of 1.0 with a weight decay of 1e-5. The pre-trainings are conducted with eight NVIDIA A100 GPUs (40GB each).
@@
To assess the DeCon framework on non scene-centric data, we pre-train SlotCon and a DeCon-SL adaptation of SlotCon with an FCN decoder on two small-scale datasets: ISIC 2017 and REFUGE described in Sec. 3.4. For ISIC dataset, we pre-train for 800 epochs with a batch size of 256 and a base learning rate of 1.0 linearly scaled with the batch size. For REFUGE, we pre-train for 2400 epochs with a batch size of 192 and the same learning rate as for ISIC. These pre-trainings are performed with one NVIDIA A6000 GPU with 48GB of memory.